A company can borrow $180 000 for 15 years. They can amortize the debt at j1 = 10%, or they can pay interest on the loan at j1 = 9% and set up a sinking fund at j1 = 7% to repay the loan. Which plan is cheaper and by how much per annum?
Answer #1

Alternative 1: Amortizing the loan:

Yearly payment required= $ 23,665.28 calculated as follows:

15 10 B C D 1 Equated Yearly Installments (1) 2 Assuming payments at the end of each period 3 Equal yearly instalment is calc

Alternative 2: Pay interest at 9% and set up sinking fund at 7% interest.

Yearly interest= Principal*Rate

Loan amount (Principal)= $180,000. Interest rate= 9% (Compounded yearly)

Therefore, yearly interest= $180,000*9%= $16,200.

Yearly payment required towards sinking fund= $7,163.03 calculated as follows:

8 9 А. B C D 3 Amount of periodical payments is calculated using the formula PMT= (PV*r)/(1-(1+r)^-n] 4 Where PMT= Periodical

Total yearly payment under alternative 2= Yearly interest at 9% interest + Yearly payment to sinking fund

=$16,200 + $7,163.03 = $23,363.03

This shows that plan 2 (Pay interest at 9% and set up sinking fund at 7% interest) is cheaper by

$23,665.28 - $23,363.03 = $302.25